# MnLargeSymbols’164 MnLargeSymbols’171

### v1

@@ Abstract
Representing vector fields by scalar or vector potentials can be a challenging task in domains with cavities or tunnels, due to the presence of harmonic fields which are both irrotational and solenoidal but may have no scalar or vector potentials. For harmonic fields normal to the boundary, which can exist in domains with cavities, it is possible to define scalar potentials with Dirichlet boundary conditions fitted to the domain’s cavities. For harmonic fields tangent to the boundary, which can exist in domains with tunnels, a similar construction was lacking. In this article we present a construction of vector potentials that yield a basis for the tangent harmonic fields. Our vector potentials are obtained by solving curl-curl problems with inhomogeneous tangent boundary conditions that are fitted to closed curves looping around the tunnels. Applied to structure-preserving finite elements, our method also provides an exact geometric parametrization of the discrete harmonic fields.
@@ 1 Introduction
Using scalar or vector potentials to represent vector fields is a powerful technique in physics and engineering. In practical computations they ensure that a vector field is either irrotational (curl-free) or solenoidal (divergence-free), which has useful applications in many branches such as astrophysics [35, 34], magnetostatics [7, 32] or fluid dynamics [28], and they are also involved in multigrid solvers [3, 23]. Beyond numerical applications, potentials are fundamental to several variational formulations, for instance in kinetic or fluid electromagnetic models [25, 27] and they play a crucial role in quantum [17] or relativistic physics [33]: a recent illustration is the elegant reformulation of the relativistic Vlasov-Maxwell equations as a wave-transport system involving Liénard-Wiechert potentials [11, 12]. Finally, their geometric properties expressed in Stokes’ theorem relate to key topological aspects of physical laws, in particular in electromagnetism [10, 19].
@@
A difficulty arises when considering domains with cavities or tunnels, where harmonic fields with vanishing traces may exist [31, 2]. Harmonic fields are unique because they are both irrotational and solenoidal, yet they may not admit scalar or vector potentials depending on their boundary conditions. Specifically, domains with cavities have harmonic fields normal to the boundary that admit scalar potentials but no vector potentials, while domains with tunnels have harmonic fields tangent to the boundary that admit vector potentials but no scalar potentials.
@@
These cavities and tunnels correspond to the 2-chain and 1-chain homology groups of the domain, whose ranks determine the dimensions of the spaces of normal and tangent harmonic vector fields, respectively [24, 2]: in particular these dimensions are topological invariants, which means that they remain unchanged under continuous deformations (with continuous inverse) of the domain. For normal harmonic fields, a construction has been proposed for scalar potentials, characterized by Laplace equations with boundary conditions fitted to the domain’s cavities [31]. For tangent harmonic fields however, a similar construction is more complex: in domains diffeomorphic to a solid torus, a common approach is to use the toroidal angle as a (piecewise smooth) scalar potential [26, 20], and this approach has been extended to a large class of domains, namely domains that can be made simply connected by removing a finite number of disjoint cuts surfaces [18, 31, 1]. A construction of proper vector potentials, however, seems to be lacking for tangent harmonic fields.
@@
In this article we propose such a construction. Our vector potentials are characterized by homogeneous curl-curl problems, with specific tangent boundary conditions fitted to closed curves that loop around the tunnels of the domain. These “tunnel curves” form a generating family for the 1-chain homology group, and the associated vector potentials yield a basis for the space of tangent harmonic vector fields. In our analysis we assume the existence of surfaces reciprocal to the tunnel curves: these surfaces are reminiscent of the cut surfaces involved in the construction of piecewise smooth scalar potentials, but here their purpose is only to establish the linear independence of the resulting harmonic fields. In particular they do not need not be disjoint, which allows us to apply our construction to domains such as hollow tori.
@@
An interesting feature of our approach is that it can also be applied in a structure-preserving finite element framework. Under some assumptions which are fulfilled by Whitney [36, 9] and higher order Nédélec elements [29, 30], and more generally by the discrete de Rham sequences constructed by Hiptmair [21] and Arnold, Falk and Winther [4, 5] in the Finite Element Exterior Calculus (FEEC) framework, a natural discretization of our construction leads to an exact geometric parametrization of the discrete harmonic fields.
@@
The article is organized as follows. In Section 2 we begin by reminding the definition of the homology groups, together with that of the harmonic spaces in the $L^{2}$ de Rham complex, and we discuss some important properties regarding their dimensions and potentials. In Section 3 we then describe our construction, which is based on a decomposition of the vector potentials in two parts: a smooth part that lifts a specific boundary condition fitted to the tunnel curves, and a correction part that guarantees that the resulting potential solves an homogeneous curl-curl problem. Section 4 is finally devoted to studying a discrete version of our construction. After listing some properties that several structure-preserving finite element frameworks possess, and discussing the case of the discrete normal harmonic fields, we show that a natural application of our approach yields a basis of the discrete tangent harmonic fields, in terms of discrete vector potentials associated to the tunnels of the domain.
@@
In our work, an important source of inspiration has been the geometric approach promoted since the 1980s by Alain Bossavit for the development of structure-preserving methods in computational electromagnetism [9, 10, 32]. For instance, a construction of discrete vector potentials for magnetic fields satisfying specific flux conditions is described in [10], within the framework of Whitney elements. We learned of Alain Bossavit’s passing as we were completing this work, which we dedicate to his memory.
@@ 2 Homology groups and harmonic function spaces
Throughout the article, we consider that $\Omega$ is an open bounded domain in $\mathbb{R}^{3}$ with a Lipschitz boundary, in the usual sense that $\partial\Omega$ is locally the graph of Lipschitz mappings, with $\Omega$ on one side only. For simplicity, we assume that $\Omega$ is connected.
@@ 2.1 Harmonic fields in the Hilbert de Rham complex
$\left\{\begin{aligned} &\mathfrak{H}^{0}=\mathbb{R},\\ &\mathfrak{H}^{1}=H(\operatorname{{\bf curl}}0;\Omega)\cap H_{0}(\operatorname{div}0;\Omega),\\ &\mathfrak{H}^{2}=H(\operatorname{div}0;\Omega)\cap H_{0}(\operatorname{{\bf curl}}0;\Omega),\\ &\mathfrak{H}^{3}=\{0\}\end{aligned}\right$. (6)
where we have used standard concise notation for the kernel spaces of the curl and divergence operators with or without boundary conditions, see e.g. [6]. If the domain is contractible (i.e., simply connected with a connected boundary), the above spaces of harmonic vector fields must vanish as a consequence of the Poincaré lemma [10]. In a general domain however this is not the case.
@@ 2.2 Singular chains and homology groups
The works of de Rham [16] have established profound connections between harmonic spaces and homology groups. Here we will follow [24] and consider singular homology groups. We remind that a singular $k$-simplex is a continuous map $\sigma:\Delta_{k}\to\Omega$ defined on the standard $k$-simplex $\Delta_{k}$, and that a (singular) $k$-chain is a formal linear combination of such singular simplices, with coefficients in $\mathbb{Z}$. Denoting by $\mathcal{C}_{k}(\Omega)$ the corresponding group of $k$-chains, and by $\partial:\mathcal{C}_{k}(\Omega)\to\mathcal{C}_{k-1}(\Omega)$ the (singular) boundary operator, we obtain a chain complex by considering the sequence
$0\xleftarrow{~~}\mathcal{C}_{0}(\Omega)\xleftarrow{~\partial~}\mathcal{C}_{1}(\Omega)\xleftarrow{~\partial~}\mathcal{C}_{2}(\Omega)\xleftarrow{~\partial~}\mathcal{C}_{3}(\Omega)\xleftarrow{~~}0.$ (7)
Here, the order from right to left corresponds to the orientation of the dual complex (2). Denoting by $\mathcal{Z}_{k}$ and $\mathcal{B}_{k}$ the $k$-cycles and $k$-boundaries, which are the kernels and ranges (in $\mathcal{C}_{k}(\Omega)$) of the boundary operator, and observing that the inclusion $\mathcal{B}_{k}\subseteq\mathcal{Z}_{k}$ follows from the sequence property $\partial\circ\partial=0$, one next defines the $k$-th homology group as the quotient
$\mathcal{H}_{k}:=\mathcal{Z}_{k}\operatorname{/}\mathcal{B}_{k}.$ (8)
Thus, an element of $\mathcal{H}_{k}$ is a class of closed $k$-chains (closed in a manifold sense, not a topological one) modulo boundary chains. Intuitively, this corresponds to the set of points, closed curves or closed surfaces (there are no closed volumes in $\Omega$) that can be obtained from each other by continuous deformations within the domain $\Omega$. An illustration is provided in Figure 1.
@@
Figure 1: Example of a non-contractible domain $\Omega$ with two cavities ($\beta_{2}=2$) and two tunnels ($\beta_{1}=2$): one that goes through the domain, and another one that is also a cavity and circles around the first tunnel. The domain boundary has three connected components which are closed surfaces (in orange, only partially drawn for the sake of visibility): $S_{0}$ is the “outer” boundary which bounds the unbounded component of $\mathbb{R}^{3}\setminus\Omega$, $S_{1}$ bounds the contractible cavity, and $S_{2}$ bounds the inner tunnel. The last two represent independent classes of the 2-chain homology group $\mathcal{H}_{2}$, while the first one is homologous to $(S_{1}+S_{2})$. Finally, the closed curves $\Gamma_{i}\subset\partial\Omega$, $i=1,2$, (in light blue) represent two independent classes of the 1-chain homology group $\mathcal{H}_{1}$. They admit reciprocal surfaces $\Sigma_{i}$ as described in Assumption 3.1, visualized by their boundary curves $\tilde{\Gamma}_{i,a}\subset\partial\Omega$, drawn in red.
The rank of the $k$-th homology group is called the $k$-th Betti number of $\Omega$, and denoted
$\beta_{k}:=\dim(\mathcal{H}_{k}).$ (9)
@@
The extremal numbers are simple: we always have $\beta_{3}=0$ since there are no closed volumes (three-dimensional cycles) in $\Omega$, and $\beta_{0}$ is the number of connected components, hence $\beta_{0}=1$ for a connected domain. The intermediate Betti numbers are more interesting and characterize the lack of contractibility of $\Omega$. In particular, $\beta_{2}$ is the number of cavities in the domain, i.e., the number of connected components of $\mathbb{R}^{3}\setminus\Omega$ which are bounded, and $\beta_{1}$ is the number of tunnels. Notice that tunnels may go through $\Omega$ like handles, or be entirely surrounded by it, which makes them cavities as well: an example is shown in Figure 1.
@@
In the sequel we will denote by $S_{i}$, $i=0,\dots,\beta_{2}$, the connected components of the boundary $\partial\Omega$ (oriented by the normal vector pointing outside of $\Omega$). They are all closed surfaces: we let $S_{0}$ be the boundary of the single unbounded connected component of $\mathbb{R}^{3}\setminus\Omega$, so that for every $i=1,\dots,\beta_{2}$, $S_{i}$ bounds a single cavity surrounded by the domain $\Omega$: these surfaces form a basis of the 2-chain homology group $\mathcal{H}_{2}$. Observe that $S_{0}$ is homologous to $(\sum_{i=1}^{\beta_{2}}S_{i})$ since the difference bounds the domain itself.
@@
Similarly, we will denote by $\Gamma_{i}$, $i=1,\dots,\beta_{1}$, a collection of oriented closed curves that generate the 1-chain homology group $\mathcal{H}_{1}$. These curves will be chosen on the boundary $\partial\Omega$, such that each one circles once around a different tunnel: a precise description will be discussed in Section 3.1 below.

### v2

@@ Abstract
Representing vector fields by potentials can be a challenging task in domains with cavities or tunnels, due to the presence of harmonic fields which are both irrotational and solenoidal but may have no scalar or vector potentials. For harmonic fields normal to the boundary, which exist in domains with cavities, the standard approach is to construct scalar potentials by solving Laplace’s equation with Dirichlet boundary conditions fitted to the closed surfaces surrounding the domain’s cavities. For harmonic fields tangent to the boundary, which exist in domains with tunnels, a similar method was lacking. In this article we present a construction of vector potentials obtained by solving curl-curl problems with inhomogeneous tangent boundary conditions fitted to closed curves looping around the tunnels. Just as the cavity surfaces represent a basis for the 2-chain homology group, these tunnel curves represent a basis for the 1-chain homology group and the corresponding vector potentials yield a basis for the tangent harmonic fields. In our analysis the linear independence of the harmonic fields is established by considering their fluxes through a collection of reciprocal surfaces. These surfaces, whose boundaries lie on the boundary of the domain and which are in intersection duality with the tunnel curves, represent a basis for the relative 2-chain homology group modulo the boundary: their existence in general domains follows from the Poincaré-Lefschetz duality. Applied to structure-preserving finite elements, our method also provides an exact geometric parametrization of the discrete harmonic fields in terms of discrete potentials.
@@ 1 Introduction
Using scalar or vector potentials to represent vector fields is a powerful technique in physics and engineering. In practical computations they ensure that a vector field is either irrotational (curl-free) or solenoidal (divergence-free), which has useful applications in many branches such as astrophysics [43, 42], magnetostatics [10, 40, 32] or fluid dynamics [36], and they are also involved in multigrid solvers [5, 29]. Beyond numerical applications, potentials are fundamental to several variational formulations, for instance in kinetic or fluid electromagnetic models [33, 35] and they play a crucial role in quantum [21] or relativistic physics [41]: a recent illustration is the elegant reformulation of the relativistic Vlasov-Maxwell equations as a wave-transport system involving Liénard-Wiechert potentials [14, 15]. Finally, their geometric properties expressed in Stokes’ theorem are linked to key topological aspects of physical laws, particularly in electromagnetism [13, 24].
@@
A difficulty arises when considering domains with cavities or tunnels, where harmonic fields with vanishing traces may exist [39, 4]. Harmonic fields are unique because they are both irrotational and solenoidal, yet they may not admit scalar or vector potentials depending on their boundary conditions. Specifically, domains with cavities have harmonic fields normal to the boundary that admit scalar potentials but no vector potentials, while domains with tunnels have harmonic fields tangent to the boundary that admit vector potentials but no scalar potentials [16].
@@
These cavities and tunnels correspond to the 2-chain and 1-chain homology groups of the domain, whose ranks (the Betti numbers) coincide with the dimensions of the spaces of normal and tangent harmonic vector fields, respectively [30, 4]: in particular these dimensions are topological invariants, which means that they remain unchanged under homeomorphisms. For normal harmonic fields, a canonical construction has been proposed where scalar potentials are characterized by Laplace problems with boundary conditions fitted to the domain’s cavities [39]. For tangent harmonic fields, a similar construction is more complex: in domains diffeomorphic to a solid torus, a common approach is to use the toroidal angle as a (piecewise smooth) scalar potential [34, 26], and this approach has been extended to domains that can be made simply connected by removing a finite number of disjoint cutting surfaces [22, 39, 3]. Such an assumption, however, imposes topological constraints which exclude simple domains such as hollow tori [9]. Moreover, a construction of proper vector potentials seemed to be lacking for tangent harmonic fields.
@@
In this article we propose such a construction. Our vector potentials are characterized by homogeneous curl-curl problems with tangent boundary conditions fitted to closed curves that loop around the tunnels of the domain. These tunnel curves form a basis family for the 1-chain homology group, and the associated vector potentials yield a basis for the space of tangent harmonic vector fields. To establish the linear independence of the resulting harmonic fields, our analysis relies on surfaces which are reciprocal to the tunnel curves. These surfaces are reminiscent of the cutting surfaces involved in the construction of piecewise smooth scalar potentials, however they are not involved in the construction step and for this reason they do not need to be disjoint or render the complementary domain simply connected. Instead, their main properties are that their own boundaries lie on the boundaries of the domain, and that they realize an intersection duality pairing with the tunnel curves. The fact that such a construction is possible for general domains is a consequence of the classical Poincaré-Lefschetz duality [11, 23].
At the discrete level, we note that several authors have developed tree-based approaches as an alternative to cutting surfaces, which have been used to compute bases of divergence-free FEM spaces in [25, 2], vector potentials of (non harmonic) prescribed vector fields in [40], and discrete loop fields in [1], which may be used to generate a basis for tangent harmonic fields.
@@
An interesting feature of our approach is that it can also be applied to compute discrete harmonic vector potentials in structure-preserving finite element spaces. Under some assumptions which are fulfilled by Whitney [44, 12] or higher order Nédélec elements [37, 38], and more generally by the discrete de Rham sequences constructed by Hiptmair [27] and Arnold, Falk and Winther [6, 7] in the Finite Element Exterior Calculus (FEEC) framework, a natural discretization of our construction leads to an exact geometric parametrization of the discrete harmonic fields.
@@
The article is organized as follows. In Section 2 we begin by reminding the definition of the homology groups, together with that of the harmonic spaces in the $L^{2}$ de Rham complex, and we discuss some important properties regarding their dimensions and potentials. In Section 3 we then discuss our main assumption and describe our construction, which is based on a decomposition of the vector potentials in two parts: a smooth part that lifts a specific boundary condition fitted to the tunnel curves, and a correction part that guarantees that the resulting potential solves an homogeneous curl-curl problem. Section 4 is finally devoted to studying a discrete version of our construction. After listing some properties that several structure-preserving finite element frameworks possess, and discussing the case of the discrete normal harmonic fields, we show that a natural application of our approach yields a basis of the discrete tangent harmonic fields, in terms of discrete vector potentials associated to the tunnels of the domain.
@@
In our work, an important source of inspiration has been the geometric approach promoted since the 1980s by Alain Bossavit for the development of structure-preserving methods in computational electromagnetism [12, 13, 40]. For instance, a construction of discrete vector potentials for magnetic fields satisfying specific flux conditions is described in [13], within the framework of Whitney elements. We learned of Alain Bossavit’s passing as we were writing this article, which we dedicate to his memory.
@@ 2 Homology groups and harmonic function spaces
Throughout the article, $\Omega$ is an open bounded domain in $\mathbb{R}^{3}$ with Lipschitz boundary, in the usual sense that $\partial\Omega$ is locally the graph of a Lipschitz mapping with $\Omega$ on one side only. For simplicity, we assume that $\Omega$ is connected.
@@ 2.1 Harmonic fields in the Hilbert de Rham complex
$\left\{\begin{aligned} &\mathfrak{H}^{0}=\mathbb{R},\\ &\mathfrak{H}^{1}=H(\operatorname{{\bf curl}}0;\Omega)\cap H_{0}(\operatorname{div}0;\Omega),\\ &\mathfrak{H}^{2}=H(\operatorname{div}0;\Omega)\cap H_{0}(\operatorname{{\bf curl}}0;\Omega),\\ &\mathfrak{H}^{3}=\{0\}.\end{aligned}\right$. (6)
Here, we have used standard concise notation for the kernel spaces of the curl and divergence operators with or without boundary conditions, see e.g. [8]. In (6) we see two spaces of harmonic vector fields: $\mathfrak{H}^{1}$ consists of fields which are tangent to the boundary (as their normal trace vanishes), whereas $\mathfrak{H}^{2}$ consists of fields which are normal to it. If the domain is contractible (i.e., simply connected with a connected boundary), the above spaces of harmonic vector fields must vanish as a consequence of the Poincaré lemma [13]. In a general domain however this is not the case.
@@ 2.2 Singular chains and homology groups
The works of de Rham [20] have established profound connections between harmonic spaces and homology groups. Here we follow [23] and consider singular homology groups in the compact domain $\overline{\Omega}$ (a topological manifold). We remind that a singular $k$-simplex (with $0\leq k\leq 3$ here) is a continuous map $\sigma:\Delta_{k}\to\overline{\Omega}$ defined on the standard $k$-simplex $\Delta_{k}$, and that a (singular) $k$-chain is a formal linear combination of such singular simplices, with coefficients in $\mathbb{Z}$. Denoting by $\mathcal{C}_{k}(\overline{\Omega})$ the corresponding group of $k$-chains, and by $\partial:\mathcal{C}_{k}(\overline{\Omega})\to\mathcal{C}_{k-1}(\overline{\Omega})$ the (singular) boundary operator, we obtain a chain complex by considering the sequence
$0\xleftarrow{~~}\mathcal{C}_{0}(\overline{\Omega})\xleftarrow{~\partial~}\mathcal{C}_{1}(\overline{\Omega})\xleftarrow{~\partial~}\mathcal{C}_{2}(\overline{\Omega})\xleftarrow{~\partial~}\mathcal{C}_{3}(\overline{\Omega})\xleftarrow{~~}0.$ (7)
Here, the order from right to left corresponds to the orientation of the dual complex (2). Denoting by $\mathcal{Z}_{k}(\overline{\Omega})$ and $\mathcal{B}_{k}(\overline{\Omega})$ the $k$-cycles and $k$-boundaries, which are the kernels and ranges (in $\mathcal{C}_{k}(\overline{\Omega})$) of the boundary operator, and observing that the inclusion $\mathcal{B}_{k}(\overline{\Omega})\subseteq\mathcal{Z}_{k}(\overline{\Omega})$ follows from the sequence property $\partial\circ\partial=0$, one next defines the $k$-th homology group as the quotient
$\mathcal{H}_{k}(\overline{\Omega}):=\mathcal{Z}_{k}(\overline{\Omega})\operatorname{/}\mathcal{B}_{k}(\overline{\Omega}).$ (8)
Thus, an element of $\mathcal{H}_{k}(\overline{\Omega})$ is a class $\sigma]=\{\sigma+\partial\tau:\tau\in\mathcal{C}_{k+1}(\overline{\Omega})\}$ of closed $k$-chains $\sigma$ (closed in a manifold sense, not a topological one), i.e. cycles, modulo boundary chains. Intuitively, this corresponds to the set of points (closed by convention), closed curves or closed surfaces that can be obtained from each other by continuous deformations within $\overline{\Omega}$. Note that in $\mathbb{R}^{3}$ no bounded volume can be closed in a manifold sense. An illustration is provided in Figure 1.
@@
Figure 1: Example of a non-contractible domain $\Omega$ with two cavities ($\beta_{2}=2$) and two tunnels ($\beta_{1}=2$): one that goes through the domain, and another one that is also a cavity and circles around the first tunnel. The domain boundary has three connected components which are closed surfaces (in orange, only partially drawn for the sake of visibility): $S_{0}$ is the “outer” boundary which bounds the unbounded component of $\mathbb{R}^{3}\setminus\Omega$, $S_{1}$ bounds the contractible cavity, and $S_{2}$ bounds the inner tunnel. The last two represent independent classes of the 2-chain homology group $\mathcal{H}_{2}(\overline{\Omega})$, while the first one is homologous to $(S_{1}+S_{2})$. Finally, the closed curves $\Gamma_{i}\subset\partial\Omega$, $i=1,2$, (in light blue) represent two independent classes of the 1-chain homology group $\mathcal{H}_{1}(\overline{\Omega})$. They admit reciprocal surfaces $\Sigma_{i}$ as described in Assumption 3.3, visualized by their boundary curves $\tilde{\Gamma}_{i,a}\subset\partial\Omega$, drawn in red.
One can verify that $\mathcal{H}_{k}(\overline{\Omega})$ is a free abelian group, akin to a vector space with integer coefficients. Its rank is called the $k$-th Betti number of $\Omega$, and denoted
$\beta_{k}:=\operatorname{rank}(\mathcal{H}_{k}(\overline{\Omega})).$ (9)
Remark 2.1.
One may also consider singular chains in the open domain $\Omega$, and the corresponding homology groups $\mathcal{H}_{k}(\Omega)$. By considering a small perturbation that moves every point of $\overline{\Omega}$ towards the interior of $\Omega$, it is not difficult to see that the two homology groups have the same ranks.
@@
The extremal numbers are simple: we always have $\beta_{3}=0$ since there are no closed volumes (three-dimensional cycles) in $\Omega$, and $\beta_{0}$ is the number of connected components, hence $\beta_{0}=1$ for a connected domain. The intermediate Betti numbers are more interesting and characterize the lack of contractibility of $\Omega$. In particular, $\beta_{2}$ is the number of cavities in the domain, i.e., the number of connected components of $\mathbb{R}^{3}\setminus\Omega$ which are bounded, and $\beta_{1}$ is the number of tunnels. Notice that tunnels may go through $\Omega$ as in handles, or be entirely surrounded by it, which makes them cavities as well: an example of both types is shown in Figure 1.
@@
In the sequel we will denote by $S_{i}$, $i=0,\dots\beta_{2}$, the connected components of the boundary $\partial\Omega$ (oriented by the normal vector pointing outside of $\Omega$). They are all closed surfaces: we let $S_{0}$ be the boundary of the single unbounded connected component of $\mathbb{R}^{3}\setminus\Omega$, so that for every $i=1,\dots\beta_{2}$, $S_{i}$ bounds a single cavity surrounded by the domain $\Omega$: these surfaces form a basis of the 2-chain homology group $\mathcal{H}_{2}$. Observe that $S_{0}$ is homologous to $(\sum_{i=1}^{\beta_{2}}S_{i})$ since the difference bounds the domain itself.
@@
Similarly, we will denote by $\Gamma_{i}$, $i=1,\dots\beta_{1}$, a family of oriented closed curves that generate the 1-chain homology group $\mathcal{H}_{1}$. These curves will be chosen on the boundary $\partial\Omega$, such that each one circles once around a different tunnel: a precise description will be discussed in Section 3.1 below.